OpenStudy (anonymous):

IM DOING K12 HELP ME METAL AND FAN

OpenStudy (anonymous):

The temperature at 8:00 a.m. was –6°F. By 4:00 p.m. the temperature was 17°F. What was the change in temperature from 8:00 a.m. to 4:00 p.m.? A. –23°F B. –11°F C. 11°F D. 23°F

OpenStudy (jadedry):

17 - (-6) = ?

OpenStudy (anonymous):

WAR

OpenStudy (anonymous):

I MEAN WAT

OpenStudy (anonymous):

11?

OpenStudy (anonymous):

A or D I did this but I forget which I think it was A

OpenStudy (jadedry):

Subtracting a negative number is the same as adding a positive one.

OpenStudy (anonymous):

uhhhh

OpenStudy (jadedry):

So 17 - (-6) = 17 + 6 = 23 = D

OpenStudy (anonymous):

so what a doing

OpenStudy (anonymous):

DDDDDDDD?????????????????

OpenStudy (jadedry):

Yes. The answer is D.

OpenStudy (anonymous):

yay

OpenStudy (anonymous):

thx